OpenStudy (anonymous):

Which of the following is a polynomial with roots 5, 4i, and −4i?

OpenStudy (anonymous):

A. f(x) = x3 − 5x2 + 20x − 16 B. f(x) = x3 − 5x2 + 16x − 80 C. f(x) = x3 − 20x2 + 5x − 16 D. f(x) = x3 − 16x2 + 80x − 5

OpenStudy (anonymous):

@mathowl609

OpenStudy (anonymous):

@Lemons

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

I apologize for the tag spam. I know that is annoying, but I really need some help.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

how does everyone always tag me?

OpenStudy (anonymous):

Your name is mathowl idk ;^; sowwy.

OpenStudy (anonymous):

its chill

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@iYuko

OpenStudy (anonymous):

Nvm I got it. It's B.

TheSmartOne (thesmartone):

Seems like I got here too late :P

OpenStudy (anonymous):

Just a bit. That's okay, though. Carry on!

imqwerty (imqwerty):

yeaB

TheSmartOne (thesmartone):

Do you want to understand how to do it? :)

OpenStudy (anonymous):

Eh not really ;-; I'm just trying to get ahead in school work before my boyfriend gets back from boot camp... Does anyone wanna help me on Google Docs with a Marine Science assignment? :D

TheSmartOne (thesmartone):

Learning it will help you more :) You need to build the foundation before you start building on top of it. But your choice :)

OpenStudy (anonymous):

Yeah I know. Algebra isn't my strong point anymore, but at this point I just need to get it done.

imqwerty (imqwerty):

jst see the product of roots

TheSmartOne (thesmartone):

^